# On the Hyperparameter Loss Landscapes of Machine Learning Algorithms

### v1

@@ Abstract
Despite the recent success in a plethora of hyperparameter optimization (HPO) methods for machine learning (ML) models, the intricate interplay between model hyperparameters (HPs) and predictive losses (a.k.a fitness), which is a key prerequisite for understanding HPO, remain notably underexplored in our community. This results in limited explainability in the HPO process, rendering a lack of human trust and difficulties in pinpointing algorithm bottlenecks. In this paper, we aim to shed light on this black box by conducting large-scale fitness landscape analysis (FLA) on $1,500$ HP loss landscapes of $6$ ML models with more than $11$M model configurations, across $67$ datasets and different levels of fidelities. We reveal the first unified, comprehensive portrait of their topographies in terms of smoothness, neutrality and modality. We also show that such properties are highly transferable across datasets and fidelities, providing fundamental evidence for the success of multi-fidelity and transfer learning methods. These findings are made possible by developing a dedicated FLA framework that incorporates a combination of visual and quantitative measures. We further demonstrate the potential of this framework by analyzing the NAS-Bench-101 landscape, and we believe it is able to faciliate fundamental understanding of a broader range of AutoML tasks.
@@ 1 Introduction
In the past decade, considerable efforts have been invested in developing hyperparameter optimization (HPO) techniques to automate the laborious task of hyperparameter (HP) tunning for machine learning (ML) models. Many successful approaches (Bergstra et al., 2011; Snoek et al., 2012; Hutter et al., 2011; Srinivas et al., 2010; Karnin et al., 2013; Li et al., 2017; Falkner et al., 2018; Awad et al., 2021) have significantly advanced this field, and they have been empirically shown to outperform both manual configuration (Hutter et al., 2019; Bischl et al., 2023; Santu et al., 2022; Yang & Shami, 2020) and random search (Bergstra & Bengio, 2012).
HPO is often casted as a black-box optimization problem (BBOP), where the goal is to search for an HP configuration $\bm{\lambda}\in\bm{\Lambda}=\Lambda_{1}\times\dots\Lambda_{n}$ with an objective value $\mathcal{L}(\bm{\lambda})$ as small as possible without any explicit knowledge of the ML loss function $\mathcal{L}:\bm{\Lambda}\to\mathbb{R}$. Existing methods (see examples above) to this end essentailly comprises $3$ key components: $i)$: a search space, $ii)$: an optimization strategy, and $iii)$: model evaluation. While the development of both efficient searching mechanisms and evaluation strategies has received considerable attention in recent years, the intricate interplay between model HPs and predictive losses, which plays a pivotal role in understanding HPO problems, remain notably underexplored. Such lack of knowledge in turn hampers the transparency and explainability (Dwivedi et al., 2023) of HPO solvers, which often function as black-boxes as well. Consequently, this results in limited human trust in HPO methods and hinders their wide-spread application (Drozdal et al., 2020; Bischl et al., 2023) Unfortunately, given the high-dimensional, hybrid nature of HP configuration space, it is far from trivial to open up such black box.
The fitness landscape metaphor, which was pioneered by Wright in 1932 in evolutionary biology, has been widely recognized as a powerful tool for analyzing BBOPs in the evolutionary computation community (Malan, 2021). It can be envisioned as a (hyper-)surface as formed by objective values, over the high-dimensional space of possible configurations (Romero et al., 2013). Since the 90s, a plethora of fitness landscape analysis (FLA) methods have been developed to conduct exploratory analysis on landscape characteristics of BBOPs (Zou et al., 2022). Such methods are useful, as they are able to extract landscape measures that are indicative of problem difficulty and how a certain search mechanism would perform on it (Smith-Miles & Lopes, 2012; Hutter et al., 2014b; Qasem & Prügel-Bennett, 2010). This knowledge would then advance the understanding of the problem characteristics (Huang & Li, 2023), assist the selection and configuration of problem solvers (Kerschke et al., 2019; Schede et al., 2022), navigate the design of new algorithms (Qasem & Prügel-Bennett, 2010), and enhance the explainability and trust for optimization (Thomson et al., 2023).
Figure 1: $2$D visualization of HP loss landscapes for: (a-d) CNN on $9$ HPs and $6,480$ configurations, (e-f) XGBoost regressor on $5$ HPs and $14,960$ configurations, under different scenarios. Colors indicate ranks of configurations (lower values are better). Coordinates are projected using UMAP.
Recently, the use of FLA in analyzing HP and the related AutoML loss landscapes has also received considerable attention. Various works have studied diverse structural characteristics of these landscapes including neutrality, modality, and fitness distance correlation (e.g., Pushak & Hoos (2022); Teixeira & Pappa (2022); Pimenta et al. (2020); Schneider et al. (2022)). However, such works suffer from limited setups and fail to interrogate the connection between landscape characteristics and the success of HP optimizers, which often run in a wide range of scenarios (e.g., different models, datasets and fidelities). It remains unclear whether the HP loss landscapes induced on different settings share certain characteristics or patterns, and how such commonalities could be potentially exploited by HP optimizers. On the other hand, we argue that existing analytical methods are insufficient to provide a comprehensive analysis and comparision on HP loss landscapes since:
  • The ability to visualize landscapes is crucial for enabling intuitive understandings of their complex topologies (Michalak, 2019). However, HP loss landscapes are notoriously difficult to visualize in a human-comprehensible fashion due to their high-dimensional nature. Some existing methods address this problem by plotting only one or two HPs each time (e.g., Friedman (2001); Akiba et al. (2019)), which fail to provide an integrated view of the global landscape structure. Other works applied dimensionality reduction techniques to project the landscape into $2$D space (e.g., Michalak (2019); Biedenkapp et al. (2018); Walter et al. (2022)), but the resulting plot is not able to preserve the overall topography as well as neighborhood structure of the landscape.
  • There is no tangible method for quantifying the similarity between different HP loss landscapes. Despite general FLA metrics could already capture informative landscape characteristics, practices in automated algorithm selection demonstrate that domain-specific metrics are also crucial as a complementary source of information for better characterizing the target problem (Smith-Miles, 2008; Smith-Miles & Lopes, 2012). However, none of the prior works have considered such perspectives when comparing HP loss landscapes.
The overarching goal of this paper is to gain an integral view of the HP loss landscapes induced on different scenarios and thereby provide new insights to the community. To this end, we develop a dedicated landscape analysis framework to enable comprehensive analysis and comparisions among HP loss landscapes. It incorporates 1: a novel neighborhood-aware HP loss landscape visualization method applicable to high-dimensions, 2: a series of FLA metrics quantifying landscape structural characteristics, and 3: $3$ similarity metrics that leverage rankings of HP configurations to allow for informative landscape similarity quantification in the HPO context. Through empirical analysis on $1,500$ landscapes across $6$ ML models and $67$ datasets with more than $11$ million configurations, we are ambitious to advance the understanding of the following four fundamental HPO scenarios:
HP Landscapes of Test Loss Versus Training Loss. ‘Overfitting’ is one of the biggest interests and concerns in the ML community (Ng, 1997; Caruana et al., 2000; Recht et al., 2019; Belkin et al., 2018; Roelofs et al., 2019; Ishida et al., 2020). However, there is a lack of in-depth understanding on how test loss correlates with training loss across a broad HP landscape, and what specific properties distinguish regions that generalize well from poorly generalized ones. In this paper, by using our fitness landscape analysis framework, we find that the test loss landscapes resemble their training counterparts in terms of both structural characteristics and performance rankings (see, e.g., Figure 1 (a) versus (b)), and configurations with small training error are likely to achieve a mild generalization error. However, significant discrepancies can also occur (see, e.g., Figure 1 (e) versus (f)) depending on both the choice of certain HP combinations and the dataset at hand. In such cases, struggling to reduce the training loss has little or even negative effect to refining the generalization loss.
HP Loss Landscapes Across Fidelities. Given the time-demanding nature of model evaluation, multi-fidelity HPO methods (Karnin et al., 2013; Kandasamy et al., 2016; Li et al., 2017; Kandasamy et al., 2017; Falkner et al., 2018; Awad et al., 2021) have achieved prominent performance by more efficient resource allocation. However, the validity of their underpinned assumption, i.e., the ranking of configuration performance would stay close to the ground truth under fidelities (Bischl et al., 2023), remains unclear (Pushak & Hoos, 2022). Our empirical results are highly inspiring to support such assumption and show that landscapes with lower fidelities are highly consistent with full-fidelity landscapes w.r.t. both structural characteristics and performance ranks (Figure 1 (c)).
HP Loss Landscapes Across Datasets. Leveraging priors obtained from previous tasks to expedite the learning process for new tasks is another promising direction for HPO (Feurer et al., 2015b; Bardenet et al., 2013; Wistuba et al., 2015b; Kim et al., 2017; Rakotoarison et al., 2022; Wistuba et al., 2015c; Vanschoren, 2018; Swersky et al., 2013). These approaches are grounded in the hypothesis that ‘knowledge’ about HPO landscapes—such as configuration quality, hyperparameter importance and their interactions (Hutter et al., 2014a; Watanabe et al., 2023b; Probst et al., 2019; van Rijn & Hutter, 2017)—is correlated across related datasets when defined under a certain distance measure. A natural question that arises is whether this knowledge remains consistent when applied to a broader range of tasks. Our results on a diverse set of $67$ datasets show that performance rankings, HP importance and their interactions, are largely shared across tasks (Figure 1 (d)).
HP Loss Landscapes Across Models. Methods rooted in Bayesian optimization (e.g., Snoek et al. (2012); Bergstra et al. (2011); Feurer et al. (2015a); Kandasamy et al. (2017)) and search space pruning (e.g., Wistuba et al. (2015a); Perrone & Shen (2019); Wang et al. (2020)) implicitly assume that the quality of configurations is locally correlated rather than a full randomness. This is also true for meta-heuristic methods (e.g., Friedrichs & Igel (2005); Lessmann et al. (2005); Cawley (2001); Guo et al. (2008)), which are even more dependent on specific landscape properties. While it may seem intuitive that HP loss landscapes would differ depending on the target ML model, in practice the fact is often that common HPO methods perform robustly for different models. This implies that, despite superficial differences, the general family of HP loss landscapes may share certain inherent patterns/properties. We verified this hypothesis by synthesizing the results from diverse FLA metrics characterizing HP loss landscape geometry combined with visual inspections (see, e.g., Figure 1 (a, e)). The results gathered from $1,500$ landscapes of $6$ ML models under different scenarios, reveal a universal picture of the HP loss landscapes. In this picture, HP loss landscapes are smooth, nearly unimodal, containing a large number of neutral regions; configurations with similar performance are locally clustered; the landscape becomes flatter around the optimum configuration.
2 HPO Landscape Analysis Methods
This section introduces our analytical framework developed to enable exploratory analysis on different HP loss landscapes and perform comparisons between them. Due to the page limit, we only provide a brief sketch of these methods here while more detailed discussions are in Appendix B.
HP Loss Landscape Construction. The HP loss landscape can be formulated as a triplet $\langle\bm{\Lambda},\mathcal{L},\mathcal{N}\rangle$ with three ingredients: $i)$ a search space $\bm{\Lambda}$ of feasible configurations that consists of pre-evaluated, discretized grid values (see Appendix F.1), $ii)$ a ML loss function $\mathcal{L}:\bm{\lambda}\to\mathbb{R}$, and $iii)$ a neighborhood structure $\mathcal{N}$ that specifies which configurations are neighbors to each other. Note that the form of $\mathcal{N}$ depends on a distance function $\mathrm{d}:\bm{\lambda}\times\bm{\lambda}\to\mathbb{N}$. Following Pushak & Hoos (2022), we define all categorical values of a HP to be distance $1$ from each other (i.e., the values are non-ordinal). For a numerical HP, we define the distance between two values to be the number of steps between them on the grid used for discretization. Such distance measure is able to mimic the tunning strategy of human experts when combined with elaborately designed grid values. Based on this, the total distance between two configurations $\bm{\lambda}_{i}$ and $\bm{\lambda}_{j}$ is then sum of the distances between the respective pairs of HP values, and we say they are neighbors to each other (i.e., $\bm{\lambda}_{j}\in\mathcal{N}(\mathcal{\bm{\lambda}}_{i})$), if $\mathrm{d}(\bm{\lambda}_{j},\bm{\lambda}_{i})=1$. Finally, the HPO landscape is constructed as a directed graph where the vertices are HP configurations and an improving edge $e_{i,j}\in\mathcal{E}$ is traced from $\bm{\lambda}_{i}$ to $\bm{\lambda}_{j}$ if $\bm{\lambda}_{j}\in\mathcal{N}(\bm{\lambda}_{i})$ and $\mathcal{L}(\bm{\lambda}_{j})<\mathcal{L}(\bm{\lambda}_{i})$. We say that a configuration $\bm{\lambda}_{\ell}$ is a local optimum if $\forall\bm{\lambda}^{\prime}\in\mathcal{N}(\bm{\lambda}_{\ell})$, we have $\mathcal{L}(\bm{\lambda}_{\ell})<\bm{\lambda}^{\prime}$. In addition, we say that $\bm{\lambda}_{j}$ is a neutral neighbor of $\bm{\lambda}_{i}$ if their performance difference is negligible ($\leq 1$‰).
Landscape Visualization. We develop a first-of-its-kind, highly interpretable method for visualizing the topography of high-dimensional HP loss landscapes by leveraging graph representation learning (Hamilton, 2020) combined with dimensionality reduction (Draganov et al., 2023) techniques. Specifically, we first extracted low-dimensional features for each node in the graph. To this end, we use HOPE (Ou et al., 2016) node embedding method because it could preserve high-order proximities between configurations. Then, we compress the obtained feature vectors into $2$ components using the UMAP (McInnes & Healy, 2018) algorithm, and thus allowing configurations to be laid out in $2$D scatter plot. To further refine the interpretability of the obtained plots, we additionally apply a linear interpolation and thereby generate a smooth landscape surface.
Table 1: Summary of the FLA metrics used in our landscape analysis framework.
Metrics
Symbol Domain What a Higher Value Implies
@@ 2 HPO Landscape Analysis Methods
$\mathcal{L}$-ast
$[-1,1]$
HP Configurations with similar $\mathcal{L}$ values are more likely to be neighbors to each other.
@@
$\rho_{a}$
$[-1,1]$
The landscape is smoother
Neutrality Distance Correlation
NDC
$[-1,1]$
The landscape is more likely to be flatter near the optimum.
@@
$\bar{\nu}$
$[0,1]$
There are many ‘plateaus’ in the landscape.
No. Local Optima
$n_{\text{lo}}$
$\mathbb{N}^{+}$
There are many ‘valleys’ or ‘peaks’ in the landscape.
Mean Basin Size
$\bar{s}_{\mathcal{B}}$
$\mathbb{R}^{+}$
The local optima are hard to be escaped from.
 ${}^{1}$ewman (2010); ${}^{2}$einberger (1990), ${}^{3}$eidys & Stadler (2001)
Quantifying Landscape Characteristics. To quantitatively assess the structural characteristics of HP loss landscapes, we employ a series of dedicated FLA metrics summarized in Table 1 as surrogate features. There are many other metrics for characterizing landscape properties (see Zou et al. (2022) for a detailed review), but our selected ones are particularly useful for this study as they cover the most essential landscape properties (i.e., modality, neutrality and smoothness) that are related to algorithm behaviors. More importantly, they are intuitive enough even for non-experts in FLA.
Landscape Similarity in Terms of Performance Ranking. The comparison of rankings of HP configurations’ performance is the essence of a large corpora of HPO methods (Hutter et al., 2019). We thereby ground our landscape similarity measure of HP loss landscapes on the consistency of their performance ranks, denoted as $R(\mathcal{L}(\bm{\lambda}))$, to allow more informative results in the HPO context. Specifically, we use $3$ statistical metrics with complementary perspectives: 1)1)1 ) Spearman’s $\rho_{s}$, it measures the association of the performance ranks of configurations in two landscapes (Spearman, 1961), 2)2)2 ) Kaggle’s Shake-up metric (Trotman, 2019), it assesses the average movement of configuration rankings across two landscapes. 3)3)3 ) The $\gamma$-set similarity (Watanabe et al., 2023a), it quantifies the ratio of overlaps between top-$10\%$ regions of two landscapes divided by their unions.
In addition to these, to investigate the consistency of HP importance and interaction under different scenarios, We apply the widely used functional ANOVA method (Hutter et al., 2014a) to assess the variance contribution of every HP $\lambda\in\bm{\lambda}$ as well as their interactions up to the $3^{\text{rd}}$ order.

### v2

# On the Hyperparameter Loss Landscapes of Machine Learning Models: An Exploratory Study

@@ Abstract
Previous efforts on hyperparameter optimization (HPO) of machine learning (ML) models predominately focus on algorithmic advances, yet little is known about the topography of the underlying hyperparameter (HP) loss landscape, which plays a fundamental role in governing the search process of HPO. While several works have conducted fitness landscape analysis (FLA) on various ML systems, they are limited to properties of isolated landscape without interrogating the potential structural similarities among them. The exploration of such similarities can provide a novel perspective for understanding the mechanism behind modern HPO methods, but has been missing, possibly due to the expensive cost of large-scale landscape construction, and the lack of effective analysis methods. In this paper, we mapped $1,500$ HP loss landscapes of $6$ representative ML models on $63$ datasets across different fidelity levels, with $11$M$+$ configurations. By conducting exploratory analysis on these landscapes with fine-grained visualizations and dedicated FLA metrics, we observed a similar landscape topography across a wide range of models, datasets, and fidelities, and shed light on several central topics in HPO.
@@ 1 Introduction
Hyperparameter optimization. In the past decade, considerable efforts have been invested in developing hyperparameter optimization (HPO) techniques to automate the laborious task of hyperparameter (HP) tunning of machine learning (ML) models, and many successful methods (e.g, [1, 2, 3, 4, 5, 6, 7, 8, 9], and see surveys [10, 11, 12, 13]) have significantly advanced this field. HPO is often cast as a black-box optimization problem, where the goal is to efficiently search for an HP configuration $\bm{\lambda}\in\bm{\Lambda}=\Lambda_{1}\times\dots\times\Lambda_{n}$ for an ML algorithm $\mathcal{A}$ that yields a loss value $\mathcal{L}(\bm{\lambda})$ as small as possible on a given dataset $\mathcal{D}$. By “black-box” it means that while we can evaluate $\mathcal{L}(\bm{\lambda})$ for any $\bm{\lambda}\in\bm{\Lambda}$, we have no access to the explicit form of the loss function $\mathcal{L}:\bm{\Lambda}\to\mathbb{R}$, nor other information like the gradients or the Hessian.
The lack of problem understanding. This lack of knowledge on $\mathcal{L}$ in turn renders HPO solvers to function as black boxes as well, and many of them are often based on rough intuition and assumptions. Albeit their prominent empirical output performance, there is a lack of understanding of the inside mechanism to support their reliability and validity. Consequently, this poses challenges for pinpointing bottlenecks of the optimizers and thus further improvement. Even worse, such lack of transparency also hinders the development of human trust in HPO, hampering its wide-spread application [11, 14]. Addressing these requires a deep understanding of the global topography of the loss function surface, which is, however, non-trivial due to its black-box nature. Also, as pointed out by a position paper [15] at ICML’24 [$\cdots$ in major ML venues, there is little incentive to engage in exploratory ML research].
Fitness landscape. The fitness landscape metaphor, which was pioneered by Sewall Wright nearly a century ago [16], is a fundamental concept in evolutionary biology [17, 18, 19, 20] and has been widely adapted for understanding black-box systems across various domains including combinatorial optimization [21, 22, 23] and physical chemistry [24, 25]. Intuitively, it can be conceived as a (hyper-)surface as formed by fitness ($\mathcal{L}$ in our context) values across a high-dimensional configuration space (Fig. 1). Each spatial location in this landscape represents a configuration (i.e., $\bm{\lambda}\in\bm{\Lambda}$), with its elevation indicating the fitness (i.e., $\mathcal{L}(\bm{\lambda})$) and optimal configurations sitting on the valleys. As an optimization process is akin to navigating downhill towards these valleys (e.g., uphill trajectories in Fig. 1), the topography of the underlying landscape contains essentially all the information needed for understanding system properties. Over the past several decades, a plethora of fitness landscape analysis (FLA) techniques [19, 26, 27] have been developed to characterize this topography, e.g., ruggedness [28], modality [29], neutrality [30], among many others.
Figure 1: Fitness landscape of an illustrative $2$D problem.
Current gaps. While several works [31, 32, 33, 34, 35] have explored various properties of the HP loss landscapes of ML models, they are often in a restricted scope (e.g., only a handful of landscapes), which limits the generalizability of their findings. More importantly, prior works often focus on individual landscapes and fail to interrogate the potential structural similarities [21] among different landscapes induced on distinct models, datasets, training fidelities, and loss functions, which are concerned by many central topics in HPO, (e.g., transfer learning [36] and multi-fidelity optimization [37]). Yet, the exploration of such similarities does not simply equate to an additive extension of prior works, but further requires both qualitative and quantitative FLA methods that could enable comprehensive comparisons among different HP loss landscapes. Unfortunately, there are no bespoke FLA metrics for this purpose, and existing landscape visualization approaches [38, 39, 40, 41, 42] struggle to tackle high-dimensionality while preserving neighborhood structure at the same time.
@@
  • HP loss landscapes on training versus test data. Overfitting is one of the biggest concerns in ML [43, 44, 45, 46, 47, 48]. While there has been empirical investigation on adaptive overffiting [45] with scattered data points from Kaggle competitions [47], there is a lack of more structured understanding, e.g., how the topography of the training loss landscape resemble that of its test counterpart?
  • HP loss landscapes with different fidelities. Given the time-demanding nature of model evaluation, multi-fidelity HPO methods [6, 49, 7, 50, 8, 9] have achieved prominent performance by more efficient resource allocation. However, while there exists an implicit assumption: the topography of HP loss landscapes with lower fidelities would stay close to the groundtruth [31, 11], it has only been indirectly verified by algorithmic performances, without landscape-level evidence.
  • HP loss landscapes across datasets. Leveraging priors obtained from previous tasks to expedite the learning process for new tasks is another promising direction for HPO [36, 51, 52, 53, 54, 55, 56, 57]. However, despite the prevalence of such transfer learning methods, there is a lack of understanding of how HP loss landscapes of the same model induced on different datasets would compare to each other?
  • HP loss landscapes across models. While it may seem intuitive that HP loss landscapes would differ depending on the target ML model, in practice the fact is often that common HPO methods could perform robustly under different ML systems. This inspires us to investigate whether the general family of HP loss landscapes may share certain inherent structural similarities.
@@
Our value proposition in this paper hinges on the hypothesis that with a deeper understanding of the underlying HP loss landscapes of ML systems, we can advance the interpretability and trustworthiness of the HPO process, and enhance the prospects of it as a tool to benefit broader scientific domains.
2 HP Loss Landscape Construction and Analysis Method
This section introduces our FLA framework developed for exploratory analysis of HP loss landscapes. We begin with a formal introduction of the HP loss landscape and how it can be constructed. We then provide a brief overview of our fine-grained landscape visualization method, selected classical FLA features, as well as our quantitative measures for landscape similarity assessment. Due to the stringent page limit, we only provide a sketch of these methods and more details are in Appendix B.
@@ 2 HP Loss Landscape Construction and Analysis Method
Landscape construction as directed graphs. Given the inherent neighborhood structure in HP loss landscapes, representing it as a graph is then a natural choice. In this model, each vertex is a configuration $\bm{\lambda}\in\bm{\Lambda}$, with the associated loss $\mathcal{L}(\bm{\lambda})$ as node attribute. Each $\bm{\lambda}^{\prime}\in\mathcal{N}(\bm{\lambda})$, is connected to $\bm{\lambda}$ via a directed edge, where the direction of each edge depends on the relative values of $\mathcal{L}(\bm{\lambda})$ and $\mathcal{L}(\bm{\lambda}^{\prime})$, always pointing to the fitter one. Such a graph-based model for the HP loss landscape inherently captures its neighborhood structure and loss dynamics. Once constructed, it then allows the implementation of many FLA methods via highly efficient graph traversal algorithms.
Landscape topography plays a crucial role in governing the search behaviors of HPO algorithms. To quantitatively characterize its features, we pick up $6$ dedicated FLA metrics, each focusing on a specific aspect, and together they offer a comprehensive view of the essential landscape properties:
  • Autocorrelation [28]: $\rho_{a}\in[-1,1]$, it measures the ruggedness of the landscape (i.e., degree of fluctuations in fitness values), where a higher value implies a smoother landscape surface.
  • Loss assortativity, $\mathcal{L}_{\text{ast}}\in[-1,1]$: adapted from the general assortativity measure of complex networks [58], this shows the tendency that configurations with similar loss values are neighbors to each other, where a large value implies more local clustering.
  • Mean neutrality, $\bar{\nu}\in[0,1]$: it is the Spearman correlation between the configurations neutralities and their distances to the global optimum; it quantifies the proportion of search space where fitness values remain constant (up to a tolerance $\epsilon=0.1\%$ in ML context) in local neighborhoods;
  • Neutrality distance correlation, $\text{NDC}\in[-1,1]$: it assesses the tendency that the terrain becomes flatter when approaching the global optimum (i.e., diminishing returns).
  • ##\##Local optima, $n_{\text{lo}}$, and mean basin size, $\bar{s}_{\mathcal{B}}$: local optima with a large basin of attraction can pose a greater challenge to optimizers, and so does a larger number of local optima.
Landscape visualization is crucial for enabling an intuitive understanding of a landscape’s topography, but is notoriously difficult because it is often in a high-dimensional space. Some prior works address this by simply plotting only one or two HP(s) each time [41, 42], which fail to capture the global topography. Others works directly applied dimensionality reduction techniques to project the landscape into a $2$D space [38, 39, 40], but the results are not able to preserve the local neighborhood structure. We tackle this by leveraging a graph embedding method, HOPE [59], to first learn a vector representation for each configuration in the landscape, which can preserve high-order proximities between nodes. We then compress the obtained embeddings into $2$D using the UMAP [60] algorithm. To further refine interpretability, we additionally apply linear interpolation to generate a continuous landscape surface. We provide more details on design choices and configurations in Appendix B.1.
[…]